# MRAG-Suite: A Diagnostic Evaluation Platform for Visual Retrieval-Augmented Generation

### v1

@@ Abstract
Multimodal Retrieval-Augmented Generation (Visual RAG) significantly advances question answering by integrating visual and textual evidence. Yet, current evaluations fail to systematically account for query difficulty and ambiguity. We propose MRAG-Suite, a diagnostic evaluation platform integrating diverse multimodal benchmarks (WebQA, Chart-RAG, Visual-RAG, MRAG-Bench). We introduce difficulty-based and ambiguity-aware filtering strategies, alongside MM-RAGChecker, a claim-level diagnostic tool. Our results demonstrate substantial accuracy reductions under difficult and ambiguous queries, highlighting prevalent hallucinations. MM-RAGChecker effectively diagnoses these issues, guiding future improvements in Visual RAG systems.
MRAG-Suite: A Diagnostic Evaluation Platform for Visual Retrieval-Augmented Generation
Yuelyu Ji University of Pittsburgh yuj49@pitt.edu
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a foundational paradigm for enhancing the factuality and reliability of large language models (LLMs) by grounding their outputs in external evidence. With recent advancements in multimodal capabilities, visual RAG systems—capable of leveraging textual and visual information simultaneously have garnered significant interest, promising substantial improvements in answering complex questions grounded in diverse modalities such as charts, plots, figures, images, and documents. Despite these advancements, current evaluation methodologies have not kept pace with these multimodal developments, primarily focusing either on purely textual contexts or closed-domain multimodal setups that provide pre-selected visual evidence. Consequently, the research community lacks rigorous evaluation benchmarks capable of systematically assessing multimodal retrieval accuracy, reasoning fidelity, and robustness to ambiguity and distractors.
Figure 1: Illustration of our multimodal claim verification process. Given a long-form answer generated by a retrieval-augmented generation (RAG) system, we extract atomic factual claims and assess whether each is supported by retrieved visual evidence. Claims are categorized as Entailed if supported by at least one image. In this example, the model correctly describes several physical traits and habitat features of a marmot , but omits the gold claim about “high-altitude environments” , which remains unsupported by the retrieved evidence.
Existing multimodal benchmarks like ScienceQA (Lu et al., 2022), MMMU (Yue et al., 2023), and MMMU-Pro (Yue et al., 2024) typically provide necessary visual contexts as part of the inputs, neglecting the retrieval component critical in real-world scenarios. Meanwhile, retrieval-centric evaluations like Visual Haystacks (Wu et al., 2024) emphasize challenging retrieval tasks but limit their scope to a single domain without fine-grained error diagnostics. Recent text-focused diagnostic frameworks such as RAGChecker and RefChekcer (Ru et al., 2024; Hu et al., 2024b) offer detailed claim-level analysis but remain restricted to textual contexts, unable to handle the additional complexity introduced by visual evidence.
To address these critical gaps, we propose MRAG-Suite, a comprehensive diagnostic evaluation platform designed explicitly for visual RAG systems. MRAG-Suite integrates eight heterogeneous datasets, spanning a wide range of real-world modalities, including charts, plots, scanned documents, slides, and open-web images. Our benchmark systematically introduces difficulty-based filtering and ambiguity-aware query variants, coupled with controlled distractor scenarios, enabling robust assessment of a model’s capability to handle ambiguity, retrieve precise evidence, and produce faithful, evidence-grounded answers.
Specifically, MRAG-Suite contributes:
@@
    Multi-domain multimodal benchmark: We consolidate and normalize datasets across diverse domains, providing unified evaluation standards and enabling direct comparisons of model performance across various multimodal tasks.
@@
    Difficulty- and ambiguity-aware filtering: By filtering trivial or easily guessable questions and introducing ambiguity through carefully constructed query variations, MRAG-Suite rigorously evaluates robustness against ambiguity and query complexity.
@@
    Claim-level multimodal diagnostic tool (MM-RAGChecker): Extending claim-level diagnostics to multimodal contexts, MM-RAGChecker systematically verifies each claim within generated answers against textual and visual evidence, enabling precise identification of hallucinations, modality biases, and retrieval accuracy.
@@ 2 Related Work
Traditional evaluation of RAG systems primarily relied on metrics like exact match and ROUGE scores, which are insufficiently fine-grained to diagnose specific failures. RAGChecker (Ru et al., 2024) introduced claim-level diagnostics for textual RAG, evaluating retrieval accuracy, factual consistency, and generation hallucinations. However, these tools remain limited to text-only contexts, lacking the capability to assess visual evidence integration and multimodal reasoning.
@@
Benchmarks such as ScienceQA (Lu et al., 2022), MMMU (Yue et al., 2023), and their advanced variants MMMU-Pro (Yue et al., 2024) evaluate models’ abilities to reason jointly over provided image and text inputs but exclude retrieval challenges. In contrast, Visual Haystacks (Wu et al., 2024) targets multimodal retrieval tasks but is constrained to single-domain visual reasoning, omitting extensive claim-level diagnostics and failing to capture multimodal reasoning intricacies across diverse domainsLi et al. (2025b, a).
@@
Recent studies reveal significant biases in multimodal RAG systems. For example, Yao et al. (Yao et al., 2025) demonstrated strong positional and modality biases, with models disproportionately relying on initial evidence items or textual over visual modalities. However, existing benchmarks have not systematically quantified these biases across multiple domains under controlled experimental conditions. MRAG-Suite specifically addresses this shortcoming by incorporating varied distractor and ambiguity scenarios to measure these biases explicitlyWang et al. (2025); Liang et al. (2025); Jiang et al. (2024); Dan et al. (2024); Lu et al. (2025); Li et al. .
@@
Prior work on multimodal QA has predominantly focused on specific domains or modalities, such as ChartMRAG (Yang et al., 2025) and VDocRAG (Tanaka et al., 2025), demonstrating the feasibility and utility of retrieval-augmented QA systems in targeted contexts. Nonetheless, these domain-specific evaluations lack a unified framework for cross-domain comparisons and fine-grained diagnostics across modalities, a gap that MRAG-Suite explicitly aims to fill.
@@ 3.2 Corpora Overview
  • WebQA Chang et al. (2022): open-web questions with both images (captions/descriptions) and short text snippets; often multi-hop and open-ended.
  • Chart-RAG Yang et al. (2025)(from ChartQA Masry et al. (2022)): chart understanding (e.g., “When did X peak?”); evidence includes the chart image plus OCR/metadata; answers are numeric/categorical values.
  • Visual-RAG Wu et al. (2025): queries that require visual knowledge (object/attribute recognition), drawn from Visual7W/FVQA/MRAG-Bench-style scenarios.
  • MRAG-Bench Hu et al. (2024a): 1,353 vision-centric MCQ converted to open-ended by taking the correct option as $a^{\text{short}}$; emphasizes fine-grained and multi-view recognition.
  • VisRAG Yu et al. (2024) subsets: Arxiv Li et al. (2024) (scholarly PDF pages), Plot Methani et al. (2019) (technical plots), Slide Tanaka et al. (2023) (slide decks), and Document Tito et al. (2022) (scanned manuals/forms). Each page is treated as a visual document with OCR text.
@@ 3.3 Normalization and Indexing
All samples are normalized to {question, short_answer, long_answer, evidence_imgs, evidence_txts}. Images use CLIP (ViT-L/14) features plus captions; text passages use dense embeddings. Visual items are indexed in FAISS; text is stored in a DPR/ElasticSearch store.
@@ 3.6 Output Specification
We evaluate along two complementary metric families rather than by answer length: (i) End-task accuracy on the short answer ($a^{\text{short}}$) via EM/Acc; and (ii) Claim-level diagnostics on the long answer ($a^{\text{long}}$) using MM-RAGChecker ; Hallucination, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, and cross-modality metrics).
@@ 4.2 Metrics
Self-knowledge measures how many entailed claims did not rely on retrieval: $\mathrm{Self\text{-}Knowledge}=\frac{\#\{\textsc{Entailment }c_{i}\ \text{with no }j(c_{i},e)=\textsc{Entailment}\ \forall e\in\mathcal{E}\}}{\#\{\textsc{Entailment }c_{i}\}}$
@@
Let $K_{\mathrm{img}}$$K_{\mathrm{txt}}$ be the numbers of retrieved images/texts. Using $s^{\mathrm{img}}_{i},s^{\mathrm{txt}}_{i}\in\{0,1\}$:
@@
Coverage of each modality:
$\mathrm{VIS\!-\!Hit@}k$ $=\frac{\#\left\{\begin{subarray}{c}\text{retrieved images that}\\ \text{entail at least one claim}\end{subarray}\right\}}{K_{\mathrm{img}}},$ (7)
$\mathrm{TXT\!-\!MissRate}$ $=1-\frac{\#\left\{\begin{subarray}{c}\text{retrieved text chunks that}\\ \text{entail any claim}\end{subarray}\right\}}{K_{\mathrm{txt}}}.$ (8)
Agreement and attribution:
$\mathrm{CMA}$ $=\frac{\sum_{i}\mathbf{1}[s^{\mathrm{img}}_{i}=1\wedge s^{\mathrm{txt}}_{i}=1]}{\sum_{i}\mathbf{1}[s^{\mathrm{img}}_{i}=1\vee s^{\mathrm{txt}}_{i}=1]},$ (9)
$\mathrm{V\!-\!HR}$ $=\frac{\#\{\,i:s^{\mathrm{img}}_{i}=1\wedge L_{i}=\textsc{Neutral}\,\}}{\#\{\,i:L_{i}=\textsc{Neutral}\,\}},$ (10)
$\mathrm{D\!-\!HR}$ $=\frac{\#\{\,i:s^{\mathrm{txt}}_{i}=1\wedge L_{i}=\textsc{Neutral}\,\}}{\#\{\,i:L_{i}=\textsc{Neutral}\,\}}.$ (11)
@@
Chart-MRAG 22.4% 19.1% 15.3% 25.7% 10.8% 11.9% 11.2% 13.5% 14.7% 29.4% 19.2% 27.6%
MRAG-Bench 32.6% 33.8% 19.5% 24.6% 28.3% 26.1% 32.4% 30.7% 38.8% 37.5% 21.9% 21.4%
VisRAG-ArXiv 55.3% 48.7% 34.6% 34.5% 35.5% 35.1% 38.6% 36.9% 43.1% 43.7% 63.9% 55.4%
VisRAG-Doc 31.1% 35.4% 20.2% 21.9% 28.5% 29.2% 27.6% 28.2% 26.9% 26.6% 39.7% 34.6%
VisRAG-Plot 36.7% 45.8% 12.3% 13.6% 35.9% 36.8% 6.4% 9.2% 51.4% 56.5% 33.5% 36.1%
VisRAG-Slide 18.5% 25.3% 8.7% 10.4% 17.6% 22.1% 24.3% 22.6% 18.2% 15.5% 20.8% 22.9%
Visual-RAG 13.8% 13.1% 9.6% 11.5% 11.7% 10.6% 8.5% 9.2% 12.1% 13.3% 11.2% 12.6%
WebQA 16.2% 14.9% 12.7% 13.4% 14.8% 17.3% 20.6% 22.1% 26.4% 27.9% 17.8% 28.7%
Table 2: EM/Accuracy performance of each model on the Filtered (retained valid samples) and Full (all samples) splits.
Dataset Model Recall Precision F1 Halluc. Faith. Self‑know. Claim R. Ctx.Prec.
@@
Table 3: Diagnostic metrics on the Filtered split (image/text pairs where applicable).
@@
Table 5: Scores (EM preferred; Accuracy used when EM is unavailable) under the three retrieval modes: DO = distractors_only, GPD = gt_plus_distractors, and GO = gt_only.
@@ 5.1 Experimental Setup
For short answers we report EM/Accuracy. For long answers we compute ROUGE-L against our reference explanations and all claim-level diagnostics using MM-RAGChecker (Hallucination Rate, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, Modality Bias). We further report the fraction of retrieved items actually used (evidence utilization). Retrieval robustness is probed under three modes: gt_only (GO), gt_plus_distractors (GPD), and distractors_only (DO).

### v2

# MRAG-Suite: A Cross-Domain Diagnostic Benchmark and Tooling for Visual Retrieval-Augmented Generation

@@ Abstract
We introduce MRAG-Suite, a cross-domain evaluation suite for Visual Retrieval-Augmented Generation (RAG) that stresses real-world retrieval, multimodal grounding, and robustness. The suite consolidates eight heterogeneous sources (web photos, charts/plots, scanned documents, slides, scholarly PDFs) under unified formats and metrics. To better reflect realistic usage, we add two-step difficulty filtering and a 200-item ambiguity split, and evaluate with three retrieval modes (gold-only, gold+ distractors, distractors-only). We also present MM-RAGChecker, a claim-level diagnostic judge for multimodal grounding, reporting hallucination, faithfulness, claim recall, context precision, and cross-modality agreement/attribution. Across open-source and proprietary VLMs, we observe large drops under difficult/ambiguous queries and in the presence of distractors, as well as strong modality biases and under-utilization of retrieved evidence. MRAG-Suite and MM-RAGChecker provide actionable, fine-grained signals for improving evidence selection and justification in visual RAG systems.
MRAG-Suite: A Cross-Domain Diagnostic Benchmark and Tooling for Visual Retrieval-Augmented Generation
Yuelyu Ji University of Pittsburgh yueluji@gmail.com          Patrick NG patrickldld@gmail.com
@@ 1 Introduction
Retrieval-augmented generation (RAG) grounds large language models (LLMs) in external evidence to improve factuality and reliability. With the rise of multimodal models, visual RAG—jointly leveraging text and images—has gained traction for complex questions involving charts, plots, figures, photos, and document pages. Yet evaluation has not kept pace: most multimodal benchmarks either assume pre-selected visuals (thus bypassing retrieval) or remain text-only, limiting our ability to measure retrieval accuracy, multimodal grounding, and robustness in realistic settings.
Existing multimodal QA datasets such as ScienceQA (Lu2022LearnTE), MMMU (Yue2023MMMUAM), and MMMU-Pro (Yue2024MMMUProAM) supply the necessary visual context in the input, side-stepping retrieval. Retrieval-centric efforts like Visual Haystacks (Wu2024VisualHA) emphasize hard look-up but focus on a single domain and lack fine-grained diagnosis. Text-focused diagnostic tools (e.g., RAGChecker, RefChecker) (ru2024ragchecker; Hu2024RefCheckerRF) provide claim-level analysis, but cannot assess how claims are grounded in visual evidence.
Figure 1: Illustration of our multimodal claim verification process. Given a long-form answer from a RAG system, we extract atomic claims and assess support from retrieved visual evidence. Claims are labeled Entailment if at least one image supports them. In this example, several traits are correctly described, but the gold claim about “high-altitude environments” remains unsupported.
To bridge these gaps, we introduce MRAG-Suite, a cross-domain diagnostic evaluation suite for visual RAG. MRAG-Suite integrates eight heterogeneous sources spanning charts/plots, natural photos, scanned documents, slides, and scholarly PDFs, and standardizes them under unified formats and metrics. To reflect real-world usage, the suite (i) applies difficulty-based filtering to de-trivialize questions, (ii) builds a 200-item ambiguity split to test underspecified queries, and (iii) evaluates three retrieval modes (gold-only, gold+ distractors, distractors-only) to probe robustness against distractors.
We further present MM-RAGChecker, a claim-level multimodal judge that verifies long-form answers against retrieved text and images. It reports hallucination and faithfulness alongside claim recall, context precision, and cross-modality behavior (e.g., agreement and attribution), yielding actionable diagnostics for error analysis and model development.
Contributions.
@@
    A unified, multi-domain visual RAG benchmark. We consolidate and normalize datasets across diverse domains, providing unified evaluation standards and enabling direct comparisons of model performance across various multimodal tasks.
@@
    Difficulty- and ambiguity-aware evaluation. Two-step filtering removes trivial items; an ambiguity subset stresses underspecified queries. Controlled distractor settings quantify robustness.
@@
    Claim-level multimodal diagnostics. MM-RAGChecker extends textual claim checking (ru2024ragchecker; Hu2024RefCheckerRF) to visual contexts, attributing support or errors to specific modalities and retrieved items.
@@ 2 Related Work
Traditional evaluation of RAG systems primarily relied on metrics like exact match and ROUGE scores, which are insufficiently fine-grained to diagnose specific failures. RAGChecker (ru2024ragchecker) introduced claim-level diagnostics for textual RAG, evaluating retrieval accuracy, factual consistency, and generation hallucinations. However, these tools remain limited to text-only contexts, lacking the capability to assess visual evidence integration and multimodal reasoning.
@@
Benchmarks such as ScienceQA (Lu2022LearnTE), MMMU (Yue2023MMMUAM), and their advanced variants MMMU-Pro (Yue2024MMMUProAM) evaluate models’ abilities to reason jointly over provided image and text inputs but exclude retrieval challenges. In contrast, Visual Haystacks (Wu2024VisualHA) targets multimodal retrieval tasks but is constrained to single-domain visual reasoning, omitting extensive claim-level diagnostics and failing to capture multimodal reasoning intricacies across diverse domainsli2025human; li2025chatmotion.
@@
Recent studies reveal significant biases in multimodal RAG systems. For example, Yao et al. (Yao2025WhoII) demonstrated strong positional and modality biases, with models disproportionately relying on initial evidence items or textual over visual modalities. However, existing benchmarks have not systematically quantified these biases across multiple domains under controlled experimental conditions. MRAG-Suite specifically addresses this shortcoming by incorporating varied distractor and ambiguity scenarios to measure these biases explicitlywang2025selfdestructivelanguagemodel; liang2025graphrag; jiang2024robustkv; dan2024evaluation; lu2025predicting; li2025frequency.
@@
Prior work on multimodal QA has predominantly focused on specific domains or modalities, such as ChartMRAG (yang2025benchmarking) and VDocRAG (Tanaka2025VDocRAGRG), demonstrating the feasibility and utility of retrieval-augmented QA systems in targeted contexts. Nonetheless, these domain-specific evaluations lack a unified framework for cross-domain comparisons and fine-grained diagnostics across modalities, a gap that MRAG-Suite explicitly aims to fill.
@@ 3.2 Corpora Overview
  • WebQA chang2022WebQA: open-web questions with both images (captions/descriptions) and short text snippets; often multi-hop and open-ended.
  • Chart-RAG yang2025benchmarking(from ChartQA masry2022chartqa): chart understanding (e.g., “When did X peak?”); evidence includes the chart image plus OCR/metadata; answers are numeric/categorical values.
  • Visual-RAG Wu2025VisualRAGBT: queries that require visual knowledge (object/attribute recognition), drawn from Visual7W/FVQA/MRAG-Bench-style scenarios.
  • MRAG-Bench hu2024mrag: 1,353 vision-centric MCQ converted to open-ended by taking the correct option as $a^{\text{short}}$; emphasizes fine-grained and multi-view recognition.
  • VisRAG Yu2024VisRAGVR subsets: Arxiv Li2024MultimodalAA (scholarly PDF pages), Plot Methani2019PlotQARO (technical plots), Slide 10.1609/aaai.v37i11.26598 (slide decks), and Document Tito2022HierarchicalMT (scanned manuals/forms). Each page is treated as a visual document with OCR text.
@@ 3.3 Normalization and Indexing
All samples are normalized to {question, short_answer, long_answer, evidence_imgs, evidence_txts}. Images use CLIP (ViT-L/14) features plus captionsradford2021learning; text passages use dense embeddings. Visual items are indexed in FAISS; text is stored in a DPR/ElasticSearch store.
@@ 3.6 Output Specification
We evaluate along two complementary metric families rather than by answer length: (i) End-task accuracy on the short answer ($a^{\text{short}}$) via EM/Acc; and (ii) Claim-level diagnostics on the long answer ($a^{\text{long}}$) using MM-RAGChecker (§4; Hallucination, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, and cross-modality metrics).
@@ 4.2 Metrics
Self-knowledge measures how many entailed claims did not rely on retrieval:
$\mathrm{Self\text{-}Knowledge}=\frac{\#\{\textsc{Entailment }c_{i}\ \text{with no }j(c_{i},e)=\textsc{Entailment}\ \forall e\in\mathcal{E}\}}{\#\{\textsc{Entailment }c_{i}\}}$
@@
We quantify cross-modality behavior with: Claim Recall (CR), Claim Precision (CP), Visual Hit@$k$ (VIS-Hit@$k$), Text Miss Rate (TXT-MissRate), Cross-Modal Agreement (CMA), Vision-attributed Hallucination Rate (V-HR), and Document/Text-attributed Hallucination Rate (D-HR). Let $\mathcal{C}$ be the set of generated atomic claims for a sample and $L_{i}\in\{\textsc{Entailment},\textsc{Contradicted},\textsc{Neutral}\}$ the verdict from MM-RAGChecker for claim $i\in\mathcal{C}$ with respect to the provided context. Let $\mathcal{I}$ and $\mathcal{T}$ denote the retrieved image and text sets; we write $v\models i$ if evidence $v$ entails claim $i$.
$s^{\mathrm{img}}_{i}$ $=\mathbf{1}\{\,\exists v\in\mathcal{I}:v\models i\,\},$ (6)
$s^{\mathrm{txt}}_{i}$ $=\mathbf{1}\{\,\exists t\in\mathcal{T}:t\models i\,\}.$ (7)
Claim-level recall/precision for $m\in\{\mathrm{img},\mathrm{txt}\}$:
[…]
$\mathrm{CP}_{m}$ $=\frac{\sum_{i\in\mathcal{C}}\mathbf{1}[s^{m}_{i}=1]}{\sum_{i\in\mathcal{C}}\mathbf{1}[\,s^{\mathrm{img}}_{i}=1\vee s^{\mathrm{txt}}_{i}=1\,]}.$ (9)
Cross-modality gaps:
@@
Coverage (restricting to top-$k$ per modality):
$\mathrm{VIS\!-\!Hit@}k$ $=\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}\!\big[\,I_{i}\vDash\exists c\,\big],$ (12)
$\mathrm{TXT\!-\!MissRate}$ $=1-\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}\!\big[\,t_{i}\vDash\exists c\,\big].$ (13)
Cross-modal agreement:
$\mathrm{CMA}=\frac{\sum_{i\in\mathcal{C}}\mathbf{1}[\,s^{\mathrm{img}}_{i}=1\wedge s^{\mathrm{txt}}_{i}=1\,]}{\sum_{i\in\mathcal{C}}\mathbf{1}[\,s^{\mathrm{img}}_{i}=1\vee s^{\mathrm{txt}}_{i}=1\,]}.$ (14)
Hallucination attribution:
$\mathrm{V\!-\!HR}$ $=\frac{\#\{\,i:\ s^{\mathrm{img}}_{i}=1\ \wedge\ L_{i}\neq\textsc{Entailment}\,\}}{\#\{\,i:\ L_{i}\neq\textsc{Entailment}\,\}},$ (15)
$\mathrm{D\!-\!HR}$ $=\frac{\#\{\,i:\ s^{\mathrm{txt}}_{i}=1\ \wedge\ L_{i}\neq\textsc{Entailment}\,\}}{\#\{\,i:\ L_{i}\neq\textsc{Entailment}\,\}}.$ (16)
Unless noted, all values are percentage points on the Filtered split, and we use $k=5$ per modality at test time.
@@
Chart-MRAG 22.4 19.1 15.3 25.7 10.8 11.9 11.2 13.5 14.7 29.4 19.2 27.6
MRAG-Bench 32.6 33.8 19.5 24.6 28.3 26.1 32.4 30.7 38.8 37.5 21.9 21.4
VisRAG-ArXiv 55.3 48.7 34.6 34.5 35.5 35.1 38.6 36.9 43.1 43.7 63.9 55.4
VisRAG-Doc 31.1 35.4 20.2 21.9 28.5 29.2 27.6 28.2 26.9 26.6 39.7 34.6
VisRAG-Plot 36.7 45.8 12.3 13.6 35.9 36.8 6.4 9.2 51.4 56.5 33.5 36.1
VisRAG-Slide 18.5 25.3 8.7 10.4 17.6 22.1 24.3 22.6 18.2 15.5 20.8 22.9
Visual-RAG 13.8 13.1 9.6 11.5 11.7 10.6 8.5 9.2 12.1 13.3 11.2 12.6
WebQA 16.2 14.9 12.7 13.4 14.8 17.3 20.6 22.1 26.4 27.9 17.8 28.7
Table 2: EM/Accuracy performance of each model on the Filtered (retained valid samples) and Full (all samples) splits. Best and Second-best are highlighted per row and split.
Dataset Model Recall Precision F1 Halluc. Faith. Self-know. Claim R. Ctx.Prec.
@@
Table 3: Diagnostic metrics on the Filtered split (image/text pairs where applicable).Best and Second-best are highlighted per row and split.All metrics are reported as percentage points except F1 (points on [0,100]).(Self-knowledge $\rightarrow$ Self-know.; Claim Recall $\rightarrow$ Claim R.; Context Precision $\rightarrow$ Ctx.Prec.)
@@
Table 5: Scores (EM preferred; Accuracy used when EM is unavailable) under the three retrieval modes: DO = distractors_only, GPD = gt_plus_distractors, GO = gt_only. † marks the best in GO; ‡ marks the best in GPD; * marks the smallest $|\mathrm{GO}-\mathrm{GPD}|$ (most robust to distractors).
@@ 5.1 Experimental Setup
For short answers we report EM/Accuracy. For long answers we compute ROUGE-L against our reference explanations and all claim-level diagnostics using MM-RAGChecker (Hallucination Rate, Faithfulness, Claim Recall, Context Precision, Self-Knowledge, Modality Bias). We further report the fraction of retrieved items actually used (evidence utilization). Retrieval robustness is probed under three modes: gt_only (GO), gt_plus_distractors (GPD), and distractors_only (DO). Unless noted, we report macro averages on the Filtered split with decoding temperature =0=0 and identical retriever outputs across models.